OpenStudy (anonymous):

Using the completing-the-square method, rewrite f(x) = x2 + 10x + 7 in vertex form

OpenStudy (anonymous):

@ganeshie8 @dan815 @Kainui @Michele_Laino @DanJS

OpenStudy (michele_laino):

hint: we can add and subtract 25, so we can write this: \[\Large \begin{gathered} f\left( x \right) = {x^2} + 10x + 7 = \hfill \\ \\ = {x^2} + 10x + 25 - 25 + 7 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

What does that do? confused

OpenStudy (michele_laino):

we have: \[\Large {x^2} + 10x + 25 = {\left( {x + 5} \right)^2}\] am I right?

OpenStudy (anonymous):

f(x)=(x+5)^2-18 f(x)=(x+10)^2+25 f(x)=(x+5)^2+7 f(x)=(x+5)^2 Those are my options btw

OpenStudy (michele_laino):

so, what is the right option? Please look at my second hint

OpenStudy (anonymous):

I'd say B

OpenStudy (michele_laino):

sorry, it is not B

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} f\left( x \right) = {x^2} + 10x + 7 = \hfill \\ \hfill \\ = {x^2} + 10x + 25 - 25 + 7 = \hfill \\ \hfill \\ = {x^2} + 10x + 25 - 18 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please keep in mind that: \[\Large {x^2} + 10x + 25 = {\left( {x + 5} \right)^2}\]

OpenStudy (anonymous):

So its A?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

Thank you!!

OpenStudy (michele_laino):

:)